# GOOD: Towards Domain Generalized Orientated Object Detection

### v1

@@ 4.2 Comparison with State-of-the-art
Method AP BF BG SV LV SP TC BC FF RA
R3Det [51] 59.4 34.1 10.2 16.6 4.0 65.5 68.7 22.4 42.5 27.8
RepPoint [26] 64.3 38.1 12.8 21.3 5.1 69.9 76.5 31.6 39.0 19.3
RoI-Tran [11] 60.6 41.0 13.6 12.3 5.6 76.0 62.3 29.3 45.3 22.2
ReDet [16] 77.9 34.6 12.5 10.6 7.6 76.8 77.1 41.7 40.4 25.2
GOOD (ReDet) 78.2 41.4 16.3 13.3 5.9 76.6 76.5 45.6 47.1 25.4
$\uparrow$0.3 $\uparrow$4.8 $\uparrow$3.8 $\uparrow$2.7 $\downarrow$1.7 $\downarrow$0.2 $\downarrow$0.6 $\uparrow$3.9 $\uparrow$6.7 $\uparrow$0.2
@@
SODA FAIR1M HRSC
RoI-Tran 0.137 0.106 0.084
GOOD (RoI-Tran) 0.091 0.058 0.043
$\downarrow$33.6$\downarrow$45.3$\downarrow$48.8%
ReDet 0.125 0.109 0.084
GOOD (ReDet) 0.086 0.049 0.038
$\downarrow$31.2$\downarrow$55.0$\downarrow$54.7%
@@
The proposed GOOD along with extensive state-of-the-art oriented object detectors are involved for experimental evaluation. For fair comparison, all the detectors are trained using the ResNet-50 backbone with a pre-trained weight on ImageNet. Besides, under each domain generalization setting, each detector is implemented under the corresponding default parameter settings with the official source code.
(a) RoI-Trans
(b) GOOD (RoI-Trans)
(c) ReDet
(d) GOOD (ReDet)
Figure 5: T-sne visualization of the cross-domain feature space. Blue, green and orange denotes the target domains of SODA, FAIR1M, HRSC, respectively. The more uniformly distributed from these domains, the better feature generalization.
@@
SODA as Source Domain The third column of Table 2 reports the results. The proposed GOOD yields an mAP of 44.65%, 47.54% and 43.88% on DOTA-1.0, -1.5 and -2.0, outperforming all the compared oriented object detectors.
DOTA-V1 as Source Domain The forth column of Table 2 reports the results. The proposed GOOD achieves the best generalization results, yielding an mAP of 57.19% and 46.21% on SODA and FAIR1M, respectively.
[…]
Generalized Ship Detection Table 4 reports the results when detecting one-type ship detection on unseen target domains. The proposed GOOD achieves the best results on all three unseen target domains, yielding an mAP of 78.97%, 48.44% and 62.08%, respectively.
4.3 Ablation Studies
On Each Component Table 5 studies the impact of each component in GOOD, namely, style hallucination, HRoI consistency learning ($\mathcal{L}_{HCL}$ in Eq. 6), rotation-aware consistency learning ($\mathcal{L}_{RCL}$ in Eq. 8) and style consistency learning ($\mathcal{L}_{SEC}$ in Eq. 10). Only implementing style hallucination slightly improves the generalization, yielding 0.99%, 0.91% and 0.57% mAP gain on DOTA-V1.0, DOTA-V1.5 and DOTA-V2.0. Implementing consistency learning on both HRoIs ($\mathcal{L}_{HCL}$) and RRoIs ($\mathcal{L}_{RCL}$) can improve the generalization ability on unseen target domains. Especially, $\mathcal{L}_{RCL}$ yields a 0.30%, 0.72% and 0.41% mAP improvement on DOTA-V1.0, DOTA-V1.5 and DOTA-V2.0 domain. Finally, the constraint on style hallucination ($\mathcal{L}_{SEC}$) can further improve the generalization.
Feature Space Visualization by t-SNE is implemented on both the proposed GOOD and its baseline model. Fig. 5 shows that the proposed GOOD allows the samples from source domain DOTA-V1 and the unseen target domain SODA, FAIR1M and HRSC to be more uniformly clustered than the baseline, which indicates its better generalization ability on unseen target domains.
[…]
On Style from Each Scale Table 7 systematically studies the impact of style hallucination on the feature representation from each scale, denoted as $\mathbf{F}^{1}$, $\mathbf{F}^{2}$, $\mathbf{F}^{3}$ and $\mathbf{F}^{4}$. The style hallucination on each scale positively contributes to the detector’s generalization ability.

### v2

@@ 4.2 Comparison with State-of-the-art
Table 3: Per-category performance comparison between the proposed GOOD and the state-of-the-art methods. FAIR1M dataset is used as the source domain. DOTA-V1.0 is used as the unseen target domain. Evaluation metric $mAP_{50}$ (in percentage %). AP: airplane; BF: baseball-field; BG: Bridge; SV: small-vehicle; LV: large-vehicle; SP: ship; TC: tennis-court; BC: basketball-court; FF: football-field; RA: roundabout. Top three results are highlighted as best, second and third, respectively.
Method AP BF BG SV LV SP TC BC FF RA mAP
R3Det Yang et al. (2021b) 59.36 34.08 10.25 16.63 3.95 65.50 68.69 22.43 42.53 27.76 35.10
RepPoint Li et al. (2022c) 64.32 38.12 12.76 21.32 5.12 69.93 76.45 31.59 39.01 19.27 37.80
RoI-Tran Ding et al. (2019) 60.64 41.04 13.63 12.34 5.58 76.02 62.32 29.32 45.32 22.23 36.82
PSCD Yu and Da (2023) 56.78 37.92 12.63 9.75 3.86 72.53 58.96 27.04 42.16 18.14 33.59
DCFL Xu et al. (2023) 75.85 39.71 13.19 12.86 4.57 75.82 71.52 34.86 44.95 23.61 39.85
ACM Xu et al. (2024) 59.31 40.87 13.08 11.69 5.28 74.68 61.75 30.17 44.63 20.39 36.17
ARC Pu et al. (2023) 58.94 38.96 12.05 12.34 4.25 73.39 58.89 27.86 43.04 19.67 34.86
JDet Pu et al. (2023) 67.93 40.15 11.87 13.07 4.86 74.21 63.06 29.78 42.68 24.05 37.05
S-DGOD Wu and Deng (2022) 50.62 33.58 10.12 8.69 2.52 63.37 54.18 22.54 36.59 13.64 29.37
DomainGen Vidit et al. (2023) 59.37 40.61 12.58 11.45 5.04 74.56 60.92 28.02 43.96 20.58 35.49
SFOD-RS Liu et al. (2024a) 57.89 38.58 11.05 10.68 4.39 74.69 60.17 27.81 42.87 20.76 34.56
PDGA Liu et al. (2024b) 60.95 40.76 12.94 12.72 5.36 74.84 61.62 28.03 43.45 21.02 36.05
ReDet Han et al. (2021b) 77.89 34.63 12.52 10.58 7.62 76.77 77.05 41.68 40.36 25.18 40.46
GOOD (Ours) 78.23 41.35 16.34 13.32 5.93 76.63 76.54 45.64 47.12 25.43 42.61
Table 4: Per-category performance comparison between the proposed GOOD and the state-of-the-art methods. FAIR1M dataset is used as the source domain. DOTA-V1.5 is used as the unseen target domain. Evaluation metric $mAP_{50}$ (in percentage %). AP: airplane; BF: baseball-field; BG: Bridge; SV: small-vehicle; LV: large-vehicle; SP: ship; TC: tennis-court; BC: basketball-court; FF: football-field; RA: roundabout. Top three results are highlighted as best, second and third, respectively.
Method AP BF BG SV LV SP TC BC FF RA mAP
R3Det Yang et al. (2021b) 59.34 34.13 10.14 27.04 3.61 64.87 68.85 23.56 41.58 14.78 34.79
RepPoint Li et al. (2022c) 63.85 37.76 12.92 32.75 4.79 67.54 74.76 30.07 40.24 3.87 36.86
RoI-Tran Ding et al. (2019) 60.62 41.07 13.43 29.82 5.82 70.31 62.06 29.37 44.36 12.89 37.02
PSCD Yu and Da (2023) 56.43 35.86 10.26 25.05 4.76 63.79 57.28 25.68 38.69 9.26 32.41
DCFL Xu et al. (2023) 72.86 40.14 14.37 28.69 5.28 74.18 66.27 31.59 44.57 12.68 39.07
ACM Xu et al. (2024) 66.04 40.25 12.86 27.47 4.65 71.23 63.58 25.86 42.04 11.54 36.58
ARC Pu et al. (2023) 64.13 36.42 12.51 25.32 4.08 65.06 61.96 28.17 41.65 11.71 35.07
JDet Xiao et al. (2024) 71.76 38.67 13.45 27.03 5.03 71.34 65.12 29.05 39.59 10.67 37.28
S-DGOD Wu and Deng (2022) 53.35 34.26 10.29 20.58 4.36 62.58 54.59 23.08 32.43 11.35 30.02
DomainGen Vidit et al. (2023) 63.88 37.58 12.16 26.36 5.15 69.74 58.73 29.47 35.62 12.26 35.08
SFOD-RS Liu et al. (2024a) 61.29 35.84 11.48 25.51 4.84 68.26 60.47 27.96 34.15 10.78 34.19
PDGA Liu et al. (2024b) 72.63 38.82 14.53 26.85 5.07 70.67 64.78 28.61 40.78 11.26 37.13
ReDet Han et al. (2021b) 77.95 37.82 12.18 29.34 7.03 77.05 77.28 39.48 41.08 11.29 41.06
GOOD (Ours) 78.04 41.54 15.93 30.58 6.14 76.82 76.43 42.33 46.42 13.13 42.70
Table 5: Per-category performance comparison between the proposed GOOD and the state-of-the-art methods. FAIR1M dataset is used as the source domain. DOTA-V2.0 is used as the unseen target domain. Evaluation metric $mAP_{50}$ (in percentage %). AP: airplane; BF: baseball-field; BG: Bridge; SV: small-vehicle; LV: large-vehicle; SP: ship; TC: tennis-court; BC: basketball-court; FF: football-field; RA: roundabout. Top three results are highlighted as best, second and third, respectively.
Method AP BF BG SV LV SP TC BC FF RA mAP
R3Det Yang et al. (2021b) 57.85 32.03 10.02 26.21 3.45 61.78 68.53 23.41 41.02 13.58 33.79
RepPoint Li et al. (2022c) 60.24 32.09 12.21 28.55 4.62 63.67 73.42 29.19 38.94 3.52 34.85
RoI-Tran Ding et al. (2019) 60.03 33.76 12.60 23.63 5.01 69.46 61.76 29.11 44.02 12.69 35.21
PSCD Yu and Da (2023) 54.82 21.69 8.45 20.58 4.39 60.71 62.35 29.68 36.16 5.37 30.86
DCFL Xu et al. (2023) 68.16 31.54 14.08 27.79 4.02 73.62 74.37 37.65 42.59 6.98 38.15
ACM Xu et al. (2024) 59.73 28.76 12.09 25.86 4.75 65.89 68.92 29.19 37.58 4.28 33.74
ARC Pu et al. (2023) 60.26 29.98 11.47 24.51 5.23 63.74 66.86 28.83 36.74 5.27 33.42
JDet Xiao et al. (2024) 64.34 32.82 13.19 21.87 5.54 69.58 69.67 31.02 41.37 6.10 35.87
S-DGOD Wu and Deng (2022) 52.65 19.54 7.87 20.73 4.18 61.53 58.96 27.89 32.63 5.09 27.86
DomainGen Vidit et al. (2023) 63.90 31.76 12.26 22.95 4.58 66.39 68.21 30.74 39.82 6.54 34.12
SFOD-RS Liu et al. (2024a) 62.67 29.85 11.72 21.27 4.35 65.37 63.49 31.58 40.95 5.67 33.74
PDGA Liu et al. (2024b) 67.57 30.51 12.93 26.84 4.97 70.16 66.28 30.13 39.69 5.84 35.22
ReDet Han et al. (2021b) 75.15 32.45 11.45 28.09 6.83 75.14 76.73 38.92 40.34 8.32 38.35
GOOD (Ours) 70.44 33.84 15.63 29.62 6.04 75.25 76.04 41.93 45.81 9.43 40.36
@@
The proposed GOOD along with extensive state-of-the-art oriented object detectors are involved for experimental evaluation. Specifically, twelve one-stage oriented object detection methods (namely, RetinaNet OBB Lin et al. (2017), FCOS OBB Tian et al. (2020), ATSS OBB Zhang et al. (2020), R3Det Yang et al. (2021b), SA2Net Han et al. (2021a), GWD Yang et al. (2021c), KLD Yang et al. (2021d), SASM Hou et al. (2022), RepPoint Li et al. (2022c), PSCD Yu and Da (2023), DCFL Xu et al. (2023), ACM Xu et al. (2024)) and seven two-stage oriented object detection methods (namely, FR OBB Ren et al. (2015), MR OBB He et al. (2017), Oriented RCNN Xie et al. (2021), RoI-Tran Ding et al. (2019), ReDet Han et al. (2021b), ARC Pu et al. (2023), JDet Xiao et al. (2024)) are included. For boarder comparison, four recent domain generalized object detectors are also involved for comparison, namely, S-DGOD Wu and Deng (2022), DomainGen Vidit et al. (2023), SFOD-RS Liu et al. (2024a) and PDGA Liu et al. (2024b). The oriented detection head is modified on S-DGOD and DomainGen, as they are not initially devised for oriented object detection.
For fair comparison, all the detectors are trained using the ResNet-50 backbone with a pre-trained weight on ImageNet. Besides, under each domain generalization setting, each detector is implemented under the corresponding default parameter settings with the official source code.
4.2.1 FAIR1M as Source Domain
The second column of Table 2 reports the results. The proposed GOOD yields an mAP of 42.61%, 42.70% and 40.36% on DOTA-1.0, -1.5 and -2.0, respectively, outperforming all the compared oriented object detectors. When embedded into ReDet Han et al. (2021b), it outperforms the ReDet baseline, which is the second best on these three unseen target domains, by 2.15%, 1.64% and 2.01% mAP, respectively. Besides, it significantly outperforms the third-best GWD Yang et al. (2021c), which only yields an mAP of 38.59%, 39.56% and 37.35% on DOTA-1.0, -1.5 and -2.0, respectively. When embedded into RoI-Tran Ding et al. (2019), it outperforms the RoI-Tran baseline by 1.12%, 0.53% and 1.10% mAP, respectively. To better understand how GOOD improves the oriented object detection performance on unseen target domains, we provide a break-down analysis on each object category.
FAIR1M $\rightarrow$ DOTA-1.0
Table 3 reports the results. The proposed GOOD outperforms the second-best by 2.15% mAP. In terms of the per-category average precision, the proposed GOOD outperforms the original ReDet on seven out of ten categories, only showing slightly inferior performance on LV, SP and TC. Finally, the proposed GOOD shows the best performance on five out of ten categories.
FAIR1M $\rightarrow$ DOTA-1.5
Table 4 reports the results. The proposed GOOD outperforms the second-best by 1.64% mAP. Besides, compared with the original ReDet, the proposed GOOD outperforms it on seven out of ten categories. Finally, the proposed GOOD shows the best performance on five out of ten categories when compared with all the runner-ups. Notably, the category-wise average precision improvement on BF, BG, BC and FF is 3.72%, 3.75%, 2.85% and 5.34%, respectively.
FAIR1M $\rightarrow$ DOTA-2.0
Table 5 reports the results. The proposed GOOD outperforms the second-best by up to 2.01% mAP. Compared with the original ReDet, the proposed GOOD is able to outperform seven out of ten categories. Further, it achieves the best performance on six out of ten categories when compared with all the runner-ups.
In summary, the proposed GOOD allows the baseline to generalize better on seven out of the ten categories, by up to 6.7% mAP on FF category, 4.8% mAP on BF category, respectively. On the other hand, a worse performance than baseline is demonstrated on three categories (LV, SP, TC) on all the three unseen target domains. It may be explained that, the density, shape and size statistic of LV, SP, TC Xia et al. (2018); Ding et al. (2021) is more varied than other categories. The proposed GOOD only constrains the orientation and content but may over-fit the shape, making it less qualified to predict the precise boxes on objects with varied unseen shapes.
4.2.2 SODA as Source Domain
The third column of Table 2 reports the results. The proposed GOOD yields an mAP of 44.65%, 47.54% and 43.88% on DOTA-1.0, -1.5 and -2.0, outperforming all the compared oriented object detectors. Specifically, when embedded into ReDet, the proposed GOOD outperforms the ReDet baseline by 2.47%, 3.29% and 2.83% mAP on these three unseen target domains, respectively. When embedded into RoI-Tran, the proposed GOOD achieves the second-best, outperforming the baseline by 2.10%, 2.67% and 1.22% mAP on these three unseen target domains, respectively.
4.2.3 DOTA-V1 as Source Domain
The forth column of Table 2 reports the results. The proposed GOOD achieves the best generalization results, yielding an mAP of 57.19% and 46.21% on SODA and FAIR1M, respectively. When embedded into ReDet, the proposed GOOD is able to outperform the baseline by 3.08% and 3.12\$ mAP on both unseen target domains. When embedded into RoI-Tran, the proposed GOOD also outperforms the baseline by 2.05% and 1.56% mAP on both unseen target domains.
4.2.4 Generalized Ship Detection
Table 6 reports the results when detecting only the ship category on unseen target domains. The proposed GOOD achieves the best results on all three unseen target domains, yielding an mAP of 78.97%, 48.44% and 62.08%, respectively. Specifically, when embedded into ReDet, the proposed GOOD outperforms the baseline by 3.08%, 2.07% and 2.37% mAP on the SODA, FAIR1M and HRSC unseen target domains, respectively. Similarly, when embedded into RoI-Tran, the proposed GOOD outperforms the baseline by 2.72%, 4.21% and 2.93% mAP on the SODA, FAIR1M and HRSC unseen target domains, respectively.
4.3 Comparison with Other Related Techniques
4.3.1 Pre-training Schemes
[…]
Method Target Domains
D-1.0 D-1.5 D-2.0
ReDet 42.18 44.25 41.05
ReDet w. CLIP 42.93 44.31 41.84
ImageNet Deng et al. (2009) 42.56 43.79 41.23
MoCo Chen et al. (2020) 42.83 44.60 41.54
BYOL Grill et al. (2020) 43.24 45.31 42.17
MAE He et al. (2022) 43.06 44.58 42.31
instance norm. Ulyanov (2016) 42.97 45.06 41.96
layer norm. Ba (2016) 42.65 44.93 42.02
GOOD (ReDet) 44.65 47.54 43.88
4.3.2 Simple Augmentation Techniques
To validate the effectiveness of the CLIP image encoder, we further compare the proposed GOOD with simpler feature normalization methods, namely, instance normalization Ulyanov (2016) and layer normalization Ba (2016). They are denoted as instance norm. and layer norm., respectively. Specifically, both methods directly implement on the source image branch, and the style hallucination has been removed. Table 7 reports the outcomes. The proposed GOOD significantly outperforms both methods, indicating its effectiveness to enrich the styles for domain generalization.
4.4 Discussion
4.4.1 Computational Cost Analysis
[…]
Table 8: Computational cost analysis between the proposed GOOD and ROI-Tran baseline. Following Ding et al. (2019), all the speed are tested on images with size of 1024$\times$1024 on a single TITAN X. #param.: parameter number (in million).
Method train speed test speed #param.
RoI-Tran 0.475 0.170 245M
GOOD (RoI-Tran) 0.728 0.170 245M
4.4.2 In-Domain Performance Analysis
It would be interesting to further study if the proposed GOOD can improve the in-domain performance of oriented object detection. The ten-category DOTA-V1.0, -V1.5 and -V2.0 (setting I in Table 1) are still used for inference, but we use the training set of DOTA-V1.0, -V1.5 and -V2.0 to train the oriented object detector. In this way, the training and inference images are within the same domain. The baseline models include both RoI-Tran Ding et al. (2019) and ReDet Han et al. (2021b). The proposed GOOD is embedded into both baselines for comparison. The results are shown in Table 9. It can be seen that, the proposed GOOD (RoI-Trans baseline) shows an mAP improvement of 2.82%, 2.65% and 2.33% on DOTA-V1.0, -V1.5 and -V2.0. Similarly, the proposed GOOD (ReDet baseline) shows an mAP improvement of 2.22%, 2.42% and 2.65% on DOTA-V1.0, -V1.5 and -V2.0. These outcomes indicate the effectiveness of GOOD on the in-domain setting.
Table 9: In-domain performance analysis between the proposed GOOD and baseline. Evaluation metric $mAP_{50}$ presented in percentage.
Method In-domain $mAP_{50}$
D-1.0 D-1.5 D-2.0
RoI-Tran 71.72 73.17 57.75
GOOD (RoI-Tran) 74.54 75.82 60.08
ReDet 78.28 79.94 59.32
GOOD (ReDet) 80.50 82.36 61.97
4.4.3 Difference from Knowledge Distillation
The proposed GOOD is a typical style hallucination based method. It is essentially different from the knowledge distillation paradigm in multiple predominant ways which prevent the data leakage. Firstly, the teacher-student structure requires the training of both the teacher and the student branch, as the teacher net needs to first learn the knowledge online. In contrast, the auxiliary branch in our method directly shares the modules and trainable parameters from the training branch on the source images. Secondly, the teacher-student structure requires the weight interaction to conduct knowledge transfer. In contrast, there is no such weight interaction between two branches in the proposed method. Thirdly, in the the teacher-student structure, the student net has to learn from the teacher’s soft prediction for knowledge transfer. In contrast, the training branch and auxiliary branch of the proposed method are both supervised by the ground truth map.
On the other hand, style hallucination is a typical paradigm to tackle domain generalization Lee et al. (2022b); Zhao et al. (2024); Yi et al. (2024a). The objective of a domain generalization model is to show well generalization to the target domains without accessing it during training. The hallucinated styles are general, not specific for the images from a certain target domain. Therefore, they are essentially treated as a kind of augmentation, rather than leak the knowledge from the target domain. Specified in our context, the CLIP image encoder definitely contains more abundant style information, as it leverages the huge amount of images from the Internet for pre-training. However, the Internet images are also general and are far different from the aerial images collected from the high-resolution satellite sensors. The outcomes in Table 7 support this point. Naively replacing the ImageNet pre-trained encoder by CLIP does not lead to a significant performance improvement on unseen target domains.
4.5 Ablation Studies
@@
D-1.0 D-1.5 D-2.0
RoI-Tran 0.116 0.078 0.130
GOOD (RoI-Tran) 0.074 0.045 0.079
$\downarrow$36.2$\downarrow$42.3$\downarrow$39.2%
ReDet 0.097 0.065 0.102
GOOD (ReDet) 0.052 0.039 0.046
$\downarrow$46.3$\downarrow$40.0$\downarrow$54.9%
4.5.1 On Each Component
Table 10 studies the impact of each component in GOOD, namely, style hallucination, HRoI consistency learning ($\mathcal{L}_{HCL}$ in Eq. 6), rotation-aware consistency learning ($\mathcal{L}_{RCL}$ in Eq. 8) and style consistency learning ($\mathcal{L}_{SEC}$ in Eq. 10).
It can be seen that, only implementing style hallucination slightly improves the generalization, yielding 0.99%, 0.91% and 0.57% mAP gain on the DOTA-V1.0, DOTA-V1.5 and DOTA-V2.0 unseen target domain, respectively. Implementing consistency learning on both HRoIs ($\mathcal{L}_{HCL}$) and RRoIs ($\mathcal{L}_{RCL}$) can improve the generalization ability on unseen target domains. Specifically, $\mathcal{L}_{HCL}$ yields a 0.07%, 0.61% and 0.46% mAP improvement on DOTA-V1.0, DOTA-V1.5 and DOTA-V2.0 unseen target domain. Similarly, $\mathcal{L}_{RCL}$ yields a 0.30%, 0.72% and 0.41% mAP improvement on DOTA-V1.0, DOTA-V1.5 and DOTA-V2.0 unseen target domain. Finally, the constraint on style hallucination ($\mathcal{L}_{SEC}$) can further improve the generalization, yielding an mAP improvement of 1.11%, 1.05% and 1.39% on the three unseen target domains, respectively.
4.5.2 Feature Space Analysis
To better understand how the proposed GOOD alleviates the domain gap for oriented object detection, we extract the feature embeddings for samples from each domain, and display them by t-SNE visualization. Both the proposed GOOD and its baseline are implemented by the above operations for comparison. Ideally, the feature embedding of a domain generalized oriented object detector is supposed to demonstrate no distribution shift for samples from different domains.
Fig. LABEL:tsne (a) and (b) compare the feature embedding distribution of the RoI-Trans baseline and the proposed GOOD. Similarly, Fig. LABEL:tsne (a) and (b) compare the feature embedding distribution of the ReDet baseline and the proposed GOOD. Both result comparisons show that the proposed GOOD allows the samples from the DOTA-V1 source domain and the unseen SODA, FAIR1M and HRSC target domains to be more uniformly clustered than the baseline, which indicates its better generalization ability on unseen target domains.
4.5.3 On Rotation Angle Prediction
@@
For more comprehensive inspection, two baselines, namely, RoI-Tran and ReDet are both involved. Table. 11 shows that when the proposed GOOD is embedded into either RoI-Tran or ReDet, the RMSD (in rad) of the rotation angle can be significantly reduced. These outcomes show the proposed GOOD can improve the rotation angle prediction on unseen target domains.
4.5.4 On Style from Each Scale
The proposed GOOD implements the style hallucination on the image encoder features of each scale. How the CLIP contributes to the feature hallucination of each scale deserves especial investigation.
Table 12 systematically studies the impact of style hallucination on the feature representation from each scale, denoted as $\mathbf{F}^{1}$, $\mathbf{F}^{2}$, $\mathbf{F}^{3}$ and $\mathbf{F}^{4}$. Specifically, we start from only hallucinating the first block feature $\mathbf{F}^{1}$, and add the block feature one by one. It is observed that, the style hallucination on each scale positively contributes to the detector’s generalization ability. Besides, the more scale features are hallucinated, the better generalization ability an oriented object detector shows.
@@ 4.5.4 On Style from Each Scale
4.5.5 On Different Distance Metric
The proposed style consistency learning ($\mathcal{L}_{SEC}$) poses the content consistency between the pre- and post- hallucinated features from an aerial image, which is implemented by minimizing the JSD distance metric. It is necessary to investigate if some other distance metrics achieve a similar effect, or whether the JSD distance can show the optimal performance on unseen target domains.